OpenStudy (anonymous):

Y^4 x^6/y^7 x simplify

OpenStudy (anonymous):

x^5/y^3

OpenStudy (solomonzelman):

\[\frac{y^4x^6}{y^7x}~~~~~~~~~~~~~correct?\] @Daredevil666, we shouldn't be givving out plain answers like what you just did

OpenStudy (anonymous):

yes that's the fraction

OpenStudy (anonymous):

Why, I just joined yesterday and I have been doing that for quite a while. Aren't we supposed to be helping them understand

OpenStudy (solomonzelman):

when you divide exponents, apply this rule. \[a^b~/~a^c~~=~a^{b-c}\]

OpenStudy (anonymous):

And giving them the answer in the process

OpenStudy (anonymous):

the how is more important than the what...

OpenStudy (anonymous):

@ShakinaMonae the answer is the first post. You basically subtract the exponents. For example, x^5=x*x*x*x*x. So I hope that simplifies it

OpenStudy (solomonzelman):

\[\frac{s^6d^4}{s^7d^2}\]think of it like this.\[\frac{s^6}{s^7} \times \frac{d^4}{d^2}\]\[\frac{s^6}{s^7}=s^{6-7}=s^{-1}=\frac{1}{s^1}=\frac{1}{s}\]\[\frac{d^4}{d^2}=d^{4-2}=d^2~~~~~~~~same~~~~as,~~\frac{d^2}{1}\]put them together and you get \[\frac{1 \times d^2}{s \times 1} ~~~~~or~~~~~\frac{d^2}{s}\]

OpenStudy (solomonzelman):

Hope this example helps you....